# Caching-Augmented Lifelong Multi-Agent Path Finding

### v1

@@ V-B Performance
Many variables can influence CAL-MAPF performance, including the number of agents, the number of caches, and input task distributions. As illustrated in Figure 4, increasing the number of caches leads to improvements in cache hit rate and makespan performance for both single and multi-port scenarios. In single-port scenario, the cache hit rate initially increases and then decreases as the number of agents grows, whereas it continues to increase in multi-port scenario. It is also observed that as the number of agents rises, CAL-MAPF performance may deteriorate compared to the baseline, even though the cache hit rate increases.
Three reasons could contribute to this scenario: (1) CAL-MAPF’s makespan performs worse than the baseline with too many agents: traffic congestions occur near the cache area and unloading port as the number of agents increases. (2) The cache hit rate initially rises with the increase in agent numbers: We use a lock mechanism to maintain items stored in the cache area. As more agents are involved, the total number of ongoing tasks increases. Since we use a shared lock for read operation and an exclusive lock for write, agents have more chances to get a read lock rather a write lock. This leads to an increase in the cache hit rate. (3) As the agent number rises, the cache hit rate decreases: Traffic congestion makes it hard for agents to complete tasks, and writing to caches becomes more challenging than reading. This nature results in caches strugglling to adapt quickly to the changes if the task distribution shifts. This second point can also clarify why CAL-MAPF outperforms the baseline as the number of caches increases: enlarging the cache area mitigates traffic congestion and enhances the cache hit rate.
@@
In Figure 5 and Figure 6, we present the performance of CAL-MAPF and the baseline across various task distributions. Figure 5 demonstrates that as the number of caches increases, CAL-MAPF’s performance improves. In the MK distribution, CAL-MAPF outperforms the baseline in most test settings. For the Real and Zhang distributions, CAL-MAPF’s performance eventually surpasses the baseline. However, with a limited number of caches or with a large number of agents, CAL-MAPF shows poor performance in both cache hit rate and makespan within the Real and Zhang distributions. One possible reason is that the low hit rate leads to agents requiring longer paths to complete a task compared to the baseline. As depicted in Figure 6, the hit rate significantly depends on the input task distribution, and the hit rate also greatly affects the final makespan performance. Traffic congestion could also contributing to CAL-MAPF’s poor performance in some test settings. Figure 3 illustrates the average frequency of agent wait actions on each grid when the number of agents reaches 256. It is evident that CAL-MAPF experiences more severe traffic congestion compared to the baseline. This also helps explain why, despite a not-so-low cache hit rate, CAL-MAPF’s performance falls below the baseline when dealing with a large number of agents, such as the performance with Real and Zhang distributions in Figure 5.
@@
A high cache hit rate and smooth traffic are crucial for the performance of CAL-MAPF. To enhance the cache hit rate, we can increase the number of caches and agents. However, both methods come with their drawbacks. The number of caches is constrained by the space available close to the unloading port. Introducing more agents may lead to severe traffic congestion, ultimately causing the hit rate to drop. Nevertheless, there are potential solutions to mitigate these adverse effects, such as implementing more efficient map designs [13] and using one-way systems near the cache and unloading ports [14]. Additionally, as we currently employ a simple task assignment policy, we could implement a more advanced policy with predictive capabilities, leveraging real warehouse task data to enhance the cache hit rate. The cache replacement policies, such as LRU and FIFO, may be overly simplistic for CAL-MAPF. Incorporating more complex policies, including learning-based approaches, could further improve the cache hit rate, especially since CAL-MAPF allows more planning time then traditional caches in computer architecture.
@@ VI Conclusion
This work presents a novel mechanism, named Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new map grid type called cache for temporary item storage and replacement. Additionally, we have devised a lock mechanism for caches to improve the stability of the planning solution. This cache mechanism was evaluated using various cache replacement policies and a spectrum of input task distributions. Through experimentation, we identified three main factors that significantly impact CAL-MAPF performance: suitable input task distribution, high cache hit rate, and smooth traffic. Overall, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions and map and agent configurations.

### v2

@@ V-B Performance
Many variables can influence the performance of CAL-MAPF, including the number of agents, the number of caches, and the distributions of input tasks. As illustrated in Figure 4, increasing the number of caches leads to improvements in cache hit rate and makespan performance for both single- and multi-port scenarios. And the cache hit rate decreases as the number of agents grows. It is also observed that as the number of agents continues to increase (from 64 to 256), there is not much difference in the makespan for both CAL-MAPF and the baseline, and the cache hit rate is dropping.
Two reasons could contribute to this scenario: (1) CAL-MAPF and baseline’s makespan remain at the same level or even worse with too many agents: traffic congestions occur in the map as the number of agents increases. (2) The CAL-MAPF cache hit rate continues to decrease as the makespan performance shows minimal differences: We employ a lock mechanism to maintain items stored in the cache area. Traffic congestion prevents many agents from reaching their target in the cache and agents always hold locks until they reach their target locations. For other agents not significantly affected by traffic congestion, they are constantly assigned new tasks and cannot obtain locks in the cache. The item distribution cannot update with input tasks, which ultimately leads to a continuous drop in the cache hit rate. This second reason can also clarify why CAL-MAPF performs better as the number of caches increases: enlarging the cache area mitigates traffic congestion and enhances the cache hit rate.
@@
In Figure 5 and Figure 6, we present the performance of CAL-MAPF and the baseline for various task distributions. Figure 5 demonstrates that as the number of caches increases, CAL-MAPF’s performance improves. In the MK, Real, and Zhang distributions, CAL-MAPF surpasses the baseline in most test settings. However, it is observable that as the number of agents increases, the improvement offered by CAL-MAPF diminishes. One possible reason is that the low hit rate leads agents to require longer paths to complete a task. As depicted in Figure 6, the hit rate significantly depends on the input task distribution, and the hit rate also greatly affects the final makespan performance.
Furthermore, in Figure 5, an abnormal increase in makespan is observed at baseline under the Zhang distribution with 256 agents. Traffic congestion could contributing to this abnormal poor performance. Figure 3 illustrates the average frequency of agent wait actions on each grid when the number of agents reaches 256 with the Zhang distribution. It is evident that both CAL-MAPF and baseline experience severe traffic congestion.
@@
High cache hit rates and smooth traffic are crucial for the performance of CAL-MAPF. To enhance the cache hit rate, we can increase the number of caches. And we can also directly add more agents to improve the makespan performance. However, both methods come with their drawbacks. The number of caches is constrained by the space available close to the unloading port. Introducing more agents can lead to severe traffic congestion, ultimately causing the hit rate to drop. Nevertheless, there are potential solutions to mitigate these adverse effects, such as implementing more efficient map designs [13] and using one-way systems near the cache and unloading ports [14]. Additionally, as we currently employ a simple task assignment policy, we could implement a more advanced policy with predictive capabilities, leveraging real warehouse task data to enhance the cache hit rate. The cache replacement policies, such as LRU and FIFO, may be overly simplistic for CAL-MAPF. Incorporating more complex policies, including learning-based approaches, could further improve the cache hit rate, especially since CAL-MAPF allows more planning time than traditional caches in computer architecture.
@@ VI Conclusion
This work presents a novel mechanism, named Caching-Augmented Lifelong MAPF (CAL-MAPF), designed to improve the performance of Lifelong MAPF. We have developed a new map grid type called cache for temporary item storage and replacement. Additionally, we have devised a locking mechanism for caches to improve the stability of the planning solution. This cache mechanism was evaluated using various cache replacement policies and a spectrum of input task distributions. Through experimentation, we identified three main factors that significantly impact CAL-MAPF performance: suitable input task distribution, high cache hit rate, and smooth traffic. In general, CAL-MAPF has demonstrated potential for performance improvements in certain task distributions and map and agent configurations.